Circular No. 7155

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V4444 SAGITTARII = NOVA SAGITTARII 1999 N. N. Samus, Sternberg Astronomical Institute, Moscow, on behalf of

IAU Commission 27, informs us that this object (cf. IAUC 7153, 7154) has been given the designation V4444 Sgr. This new policy of rapidly assigning permanent designations for occasional unusual Milky Way variable stars that are announced on these Circulars was instituted following discussions with the Central Bureau.

W. Liller, Viña del Mar, Chile, reports that another objective-prism spectrogram (cf. IAUC 7153) secured on Apr. 29.208 UT shows the Hα emission much brighter than on Apr. 27 (it is now 2.1 times brighter than the neighboring continuum), and Hβ emission has become visible. Additional broadband V CCD magnitudes: Apr. 28.152 UT, 8.09; 29.219, 8.719.

CCD photometry by D. Hanžl, Brno, Czech Republic (0.40-m reflector; comparison star GSC 6850.1335, V = 8.59, B − V = −0.018): Apr. 30.108 UT, V = 9.04 ± 0.03, B − V = +0.07 ± 0.03.

Visual magnitude estimates: Apr. 27.457 UT, 7.7 (A. Jones, Nelson, New Zealand); 27.699, 7.7 (Jones); 27.751, 7.7 (Jones); 28.073, 8.5 (P. Schmeer, Bischmisheim, Germany); 28.331, 8.2 (J. G. de S. Aguiar, Campinas, Brazil); 29.097, 9.0 (K. Hornoch, Lelekovice, Czech Republic); 29.105, 8.8 (Schmeer); 29.636, 9.0 (A. Pearce, Nedlands, W. Australia).

V4334 SAGITTARII G. Jacoby, Kitt Peak National Observatory (KPNO), reports: “The

unusual variable star V4334 Sgr (cf. IAUC 6322) has continued to fade and become redder (cf. IAUC 7095). Photometric measures obtained by R. Ciardullo and J. Feldmeier, Pennsylvania State University, at the KPNO 4-m telescope in B and V on Mar. 21.51 UT and in R and I on Mar. 22.52 yield: V = 20.08, B − V = +3.17, V − R = +2.75, and V − I = +4.99. Observations by K. Hinkle and R. Joyce at the KPNO 2.1-m telescope on Apr. 8.42 at 1083 and 2230 nm yield V −m(1083) = +8.3 and V −m(2230) = +13.5, assuming that V4334 Sgr had not faded further between Mar. 21 and Apr. 8. All photometry has an accuracy of < 0.2 mag. If the infrared emission is from a blackbody, then the m(1083-nm)–m(2230-nm) color suggests that its temperature is ∼ 1285 K. The observed optical fluxes, however, are not reproduced accurately by the prediction of a pure 1285-K blackbody, being brighter at B, V, and R relative to the prediction. Emission in the region near 1083 nm appeared to be continuous.”

1999 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green